# Graph Neural Networks for the Offline Nanosatellite Task Scheduling Problem

### v3

@@ 3.1 MILP Formulation
$\phi_{j,t}\geq x_{j,t},$ $\leavevmode\nobreak\ \forall j\in\mathcal{J},\,t=1$ (2a)
$\phi_{j,t}\geq x_{j,t}-x_{j,(t-1)},$ $\leavevmode\nobreak\ \forall j\in\mathcal{J},\,\forall t\in\mathcal{T}:t>1$ (2b)
$\phi_{j,t}\leq x_{j,t},$ $\leavevmode\nobreak\ \forall j\in\mathcal{J},\,\forall t\in\mathcal{T}$ (2c)
$\phi_{j,t}\leq 2-x_{j,t}-x_{j,(t-1)},$ $\leavevmode\nobreak\ \forall j\in\mathcal{J},\,\forall t\in\mathcal{T}:t>1$ (2d)
@@
$\phi_{j,t}\in\{0,1\},$ $\leavevmode\nobreak\ \forall j\in\mathcal{J},t\in\mathcal{T}$ (2n)
$x_{j,t}\in\{0,1\},$ $\leavevmode\nobreak\ \forall j\in\mathcal{J},t\in\mathcal{T}.$ (2o)
@@
are used, which take a positive value if, and only if, job $j$ was not running at time $t-1$, but started running at time $t$. Constraints (2a) to (2d) are used to ensure this behavior. We also consider that jobs may run only during a time window defined by parameters $w^{\min}_{j}$ and $w^{\max}_{j}$, which is ensured by constraints (2e) and (2f). Such behavior is necessary to ensure that a payload, for instance, runs only when passing above a certain territory.
@@
$\sum_{j=1}^{J}q_{j}x_{j,t}\leq r_{t}+\gamma\leavevmode\nobreak\ V_{b},$ $\forall t\in\mathcal{T}$ (3a)
@@
$\text{SoC}_{t+1}=\text{SoC}_{t}+\frac{i_{t}\leavevmode\nobreak\ e}{60\leavevmode\nobreak\ Q},$ $\forall t\in\mathcal{T}$ (3d)
@@ 4.1 MILP Problems as Inputs for GNNs
$\boldsymbol{c}=\begin{bmatrix}1\\ 2\\ 3\end{bmatrix};\leavevmode\nobreak\ \\ A=\begin{bmatrix}1&2&0\\ 0&1&-1\\ 5&0&1\end{bmatrix};\leavevmode\nobreak\ \\ \boldsymbol{b}=\begin{bmatrix}2\\ 1\\ 4\end{bmatrix}.$ (11)
@@
[…]
Figure 3: Bipartith graph representation of an MILP with three variables ($x_{1}$, $x_{2}$, and $x_{3}$), three constraints ($C_{1}$, $C_{2}$, and $C_{3}$), and $A$ as in Eq. (11).
@@ 5.2 SatGNN
We name SatGNN the base model proposed for all experiments. An overview of the components of our model can be seen in the diagram of Figure 4. We follow the approach of embedding the ONTS problem instances as bipartite graphs and feeding them to a GNN (see Sec. 4). Because of the convolutional nature of the message-passing iterations of the GNNs, the model can deal with arbitrary-sized graphs, which enables the same GNN to handle optimization problems with varying numbers of variables and constraints. In other words, the GNN can handle ONTS instances with varying numbers of jobs, which is useful, e.g., for iterating over the number of jobs during the mission design to maximize QoS.
The input graph is enhanced with feature vectors for every node. Let $G=(V,E,w)$ be the graph representation of an instance $I\in\mathcal{I}$ as described in Sec. 4.1, with $V=V_{\textrm{var}}\cup V_{\textrm{con}}$ the set of nodes and $w:E\longrightarrow\mathbb{R}$ the edge weight function based on the variables coefficients on the constraints. We name $\boldsymbol{f}_{v}$ the feature vector associated with node $v$. We add four features to constraint nodes and six to variable nodes, as described in Table 2.
@@
${\rm NN}_{\rm var}:\mathbb{R}^{6}$ +$\longrightarrow\mathbb{R}^{d}_{+}$
@@
${\rm NN}_{\rm con}:\mathbb{R}^{4}$ +$\longrightarrow\mathbb{R}^{d}_{+}$
@@
Figure 4: SatGNN architecture overview, where $G$ is the bipartite graph representation of an MILP problem instance, $H$ variables represent the sets of hidden features of the nodes, and $NN_{\cdot}$ are neural networks. The connection of $G$ and the $GraphConv$ operators represents both the weights of the edges and the neighborhood information.
@@ 5.3.1 Feasibility Classification
Given an instance of the ONTS problem $I\in\mathcal{I}$ and a candidate solution $\boldsymbol{\hat{z}}$, we want the model to determine whether $\hat{\boldsymbol{z}}$ is a feasible solution or not, that is, $p(\hat{\boldsymbol{z}}\in Z|I)$, where $Z\subset\{0,1\}^{2JT}$ is the set of feasible solutions. The candidate solution is fed to the model as features, adding an extra dimension to $\boldsymbol{f}_{v_{\rm var}}$. As briefly described in Section 5.2, the feasibility classification is a graph classification task, thus, we aggregate the output of SatGNN before the sigmoid function, i.e., the outputs of ${\rm NN}_{\rm out}$ for each variable node are aggregated to generate the predicted logit, before applying the sigmoid to have the predicted probability $\hat{p}(\hat{\boldsymbol{z}}\in Z|I)$. More precisely, we can write
$\hat{p}(\hat{\boldsymbol{z}}\in Z|I)=\frac{1}{|V_{\rm var}|}\sum_{v_{\rm var}\in V_{\rm var}}{\rm NN}_{\rm out}\left(\boldsymbol{h}^{(L)}_{v_{\rm var}}\right),$
where $\boldsymbol{h}^{(L)}_{v_{\rm var}}$ is computed as described in Sec. 5.2, and $V_{\rm var}$ is the set of variable nodes, as described in Sec. 4.1. The model is trained to minimize the binary cross-entropy (BCE) between the predicted probability and the actual probability
$\min_{\theta}\mathcal{L}_{\rm feas}(\theta)=\sum_{(I,\hat{\boldsymbol{z}},y)\in\mathcal{D}_{\rm feas}}{\rm BCE}({\rm SatGNN}(I,\hat{\boldsymbol{z}};\theta),y),$
where $y\in\{0,1\}$ indicates whether $\hat{\boldsymbol{z}}$ is feasible ($y=1$) or not ($y=0$), and $\theta$ encapsulates all trainable parameters of SatGNN, which includes the parameters of the graph convolutions and the parameters of ${\rm NN}_{\rm con}$, ${\rm NN}_{\rm var}$, and ${\rm NN}_{\rm out}$.
The training data is built with instances from $\mathcal{D}$ (see Sec. 5.1). For each instance $I$ such that $(I,Z^{\star})\in\mathcal{D}$, beyond $Z^{\star}$ as a set of feasible solutions, we also associate with the instance two other sets of candidate solutions $Z_{R},Z_{N}\subset Z$, such that the dataset can be described as
$\mathcal{D}_{\rm feas}=\left\{(I,\hat{\boldsymbol{z}},y)\in\mathcal{I}\times Z\times\{0,1\}:(I,Z^{\star})\in\mathcal{D},\hat{\boldsymbol{z}}\in Z^{\star}\cup Z_{R}\cup Z_{N},y=p(\hat{\boldsymbol{z}}\in Z|I)\right\}.$
Note that $p(\hat{\boldsymbol{z}}\in Z|I)\in\{0,1\}$, as it is computed analytically by evaluating the constraints.
The first set is built with randomly sampled instances
$Z_{R}=\left\{\boldsymbol{\hat{z}}\in Z:\boldsymbol{\hat{z}}\sim\mathcal{U}\left[\{0,1\}^{JT}\right]\right\},$
where $\mathcal{U}\left[\{0,1\}^{JT}\right]$ denotes a uniform distribution over the set of all possible assignments for the binary variables. The second set is built with candidate solutions from the neighborhood of the solutions in $Z^{\star}$,
$Z_{N}=\left\{\boldsymbol{\hat{z}}\in Z:\|\boldsymbol{\hat{z}}-\boldsymbol{z}^{\star}\|_{0}\leq\eta,\boldsymbol{z}^{\star}\in Z^{\star}\right\},$
where $\eta\in\mathbb{N}$ limits the size of the neighborhood. In our experiments, we build candidate solutions in $Z_{N}$ by randomly flipping at most $\eta$ values from a randomly selected candidate solution from $Z^{\star}$. Intuitively, $Z^{\star}$ and $Z_{N}$ are used to teach the model the limits of the feasible region, while $Z_{R}$ carries general information about the domain of the problem.
@@ 5.3.2 Optimal Solution Prediction
We want to train the SatGNN model to predict the variables’ bias in optimal solutions for instances of the ONTS problem $I\in\mathcal{I}$. More precisely, we want the model to estimate the probability that each binary variable takes a positive value in the optimal solution, which we will denote as
$\boldsymbol{\hat{p}}(\boldsymbol{z}^{\star}=1|I)=\begin{bmatrix}\hat{p}(z^{\star}_{1,1}=1|I)\\ \vdots\\ \hat{p}(z^{\star}_{J,T}=1|I)]\end{bmatrix}.$
Therefore, we have a node classification task for which we apply the sigmoid function at the output of SatGNN for each variable node such that
${\rm SatGNN}:\mathcal{I}\times\Theta$ $\longrightarrow[0,1]^{JT}$
$I;\theta$ $\longmapsto\hat{\boldsymbol{p}}(\boldsymbol{z}^{\star}=1|I)={\rm SatGNN}(I;\theta),$
where $\theta\in\Theta$ is the vector of trainable parameters. Note that the most probable candidate solution $\boldsymbol{\hat{z}}=(\hat{\boldsymbol{x}},\hat{\boldsymbol{\phi}})\in\{0,1\}^{2JT}$ is easily obtained given the model’s prediction as
$\hat{x}_{j,t}=\lceil\hat{p}(x^{\star}_{j,t}=1|I)\rfloor,\forall(j,t)\in\mathcal{J}\times\mathcal{T}$
and similarly for $\hat{\phi}_{j,t}$.
One way to train the model is to minimize the BCE between the predicted probability and the single best solution known for each instance, minimizing the loss
$\min_{\theta}\mathcal{L}_{\rm opt\text{-}b}(\theta)=\sum_{(I,\boldsymbol{z}^{\star})\in\mathcal{D}_{\rm opt\text{-}b}}{\rm BCE}({\rm SatGNN}(I;\theta),\boldsymbol{z}^{\star}),$
where the dataset is built such that $\boldsymbol{z}^{\star}$ is the best solution available for each instance in $\mathcal{D}$, i.e.,
$\mathcal{D}_{\rm opt\text{-}b}=\left\{(I,\boldsymbol{z}^{\star})\in\mathcal{I}\times Z:(I,Z^{\star})\in\mathcal{D},\boldsymbol{z}^{\star}=\arg\max_{\hat{\boldsymbol{z}}\in Z^{\star}}{\rm QoS}(\hat{\boldsymbol{z}};\boldsymbol{u}_{I})\right\},$
@@
Another way is through the approach proposed by Nair et al. (2020), in which multiple solutions are used as targets, being weighted by their objective value. Note that the best-known solution is also used in this approach, along with sub-optimal solutions. In this approach, the optimization can be described as
$\min_{\theta}\mathcal{L}_{\rm opt\text{-}m}(\theta)=\sum_{(I,Z^{\star})\in\mathcal{D}_{\rm opt\text{-}m}}\sum_{\boldsymbol{z}^{\star}\in Z^{\star}}w(\boldsymbol{z}^{\star};Z^{\star},I)\cdot{\rm BCE}({\rm SatGNN}(I;\theta),\boldsymbol{z}^{\star}),$
where
$w(\boldsymbol{z}^{\star};Z^{\star},I)=\frac{\exp({\rm QoS}(\boldsymbol{z}^{\star};\boldsymbol{u}_{I}))}{\sum_{\boldsymbol{z}\in Z^{\star}}\exp({\rm QoS}(\boldsymbol{z};\boldsymbol{u}_{I}))}$
is the weight associated with each candidate solution found by the solver, which is built such that the better the objective, the higher the weight, and, for each instance, the weights add up to 1. Note that we can directly take $\mathcal{D}_{\rm opt\text{-}m}\subseteq\mathcal{D}$.
@@ 5.4 SatGNN-based Heuristic
The output of the SatGNN as trained for the optimal solution prediction, that is, the predicted variables’ bias, can be rounded to the closest integer (either 0 or 1) to generate the most probable candidate solution. However, this is not expected to perform well as we have no feasibility guarantees; thus, any error in the network may generate an infeasible solution. Given that ONTS problems of moderate size already have thousands of variables [^3] The smallest instance evaluated by Rigo et al. (2022) has 1746 binary variables, considering both $x$ and $\phi$ variables., infeasibility is likely to occur. Therefore, we propose using the deep learning model to aid the optimization in three approaches: warm-starting, early-fixing, and trust region.
@@ 5.4.1 Warm-starting
Another straightforward approach to using the information learned by the model is to provide (partial) solutions to a solver, warm-starting the optimization. For example, the SCIP solver (Bestuzheva et al., 2021) accepts complete and partial solutions to the problem, which are used to guide the inner heuristics during the optimization process. We use the output of the model to determine which variables will compose the partial solution provided to the solver based on the confidence of the model’s prediction. In other words, the closer the model’s probability output is to the extreme values (0 or 1), the more confident it is of that assignment.
Formally, let $\hat{\boldsymbol{p}}(\boldsymbol{z}^{\star}=1|I)$ be the output of SatGNN given an instance $I\in\mathcal{I}$ as defined in Sec. 5.3.2. Then, we will denote $\boldsymbol{\kappa}(\boldsymbol{z}^{\star}|I)\in[0,1]^{2JT}$ the confidence the model has in each binary variable. For variable $x_{j,t}$ of instance $I$, given that the model’s output is $\hat{p}(x_{j,t}^{\star}=1|I)$, then we can describe the respective confidence $\kappa(x_{j,t}|I)$ as
$\kappa(x_{j,t}|I)=\begin{cases}\hat{p}(x_{j,t}^{\star}=1|I),&\text{if}\quad\hat{p}(x_{j,t}^{\star}=1|I)\geq 0.5\\ \hat{p}(x_{j,t}^{\star}=0|I)=1-\hat{p}(x_{j,t}^{\star}=1|I),&\text{otherwise}\end{cases}.$
The confidence associated with the $\phi_{j,t}$ variables can be described similarly. Note that the confidence is always between 0.5 and 1.
@@
Note that warm-starting does not configure a heuristic solution to the problem, as it just modifies the behavior of the heuristics already present in the algorithmic solution. In other words, warm-starting will only modify the order of exploring the B&B tree. Therefore, we maintain optimality and feasibility guarantees, even in the worst-case scenario when the model provides the solver with bad, infeasible solutions.

### v4

@@ 3.1 MILP Formulation
$\phi_{j,t}\geq x_{j,t},$ $\penalty 10000\ \forall j\in\mathcal{J},\,t=1$ (2a)
$\phi_{j,t}\geq x_{j,t}-x_{j,(t-1)},$ $\penalty 10000\ \forall j\in\mathcal{J},\,\forall t\in\mathcal{T}:t>1$ (2b)
$\phi_{j,t}\leq x_{j,t},$ $\penalty 10000\ \forall j\in\mathcal{J},\,\forall t\in\mathcal{T}$ (2c)
$\phi_{j,t}\leq 2-x_{j,t}-x_{j,(t-1)},$ $\penalty 10000\ \forall j\in\mathcal{J},\,\forall t\in\mathcal{T}:t>1$ (2d)
@@
$\phi_{j,t}\in\{0,1\},$ $\penalty 10000\ \forall j\in\mathcal{J},t\in\mathcal{T}$ (2n)
$x_{j,t}\in\{0,1\},$ $\penalty 10000\ \forall j\in\mathcal{J},t\in\mathcal{T}.$ (2o)
@@
are used, which take a positive value if, and only if, job $j$ was not running at time t1t-1, but started running at time $t$. Constraints (2a) to (2d) are used to ensure this behavior. We also consider that jobs may run only during a time window defined by parameters $w^{\min}_{j}$ and $w^{\max}_{j}$, which is ensured by constraints (2e) and (2f). Such behavior is necessary to ensure that a payload, for instance, runs only when passing above a certain territory.
@@
$\sum_{j=1}^{J}q_{j}x_{j,t}\leq r_{t}+\gamma\penalty 10000\ V_{b},$ $\forall t\in\mathcal{T}$ (3a)
@@
$\text{SoC}_{t+1}=\text{SoC}_{t}+\frac{i_{t}\penalty 10000\ e}{60\penalty 10000\ Q},$ $\forall t\in\mathcal{T}$ (3d)
@@ 4.1 MILP Problems as Inputs for GNNs
$\boldsymbol{c}=\begin{bmatrix}1\\ 2\\ 3\end{bmatrix};\penalty 10000\ \\ A=\begin{bmatrix}1&2&0\\ 0&1&-1\\ 5&0&1\end{bmatrix};\penalty 10000\ \\ \boldsymbol{b}=\begin{bmatrix}2\\ 1\\ 4\end{bmatrix}$. (11)
@@
$x_{1}$$x_{2}$$x_{3}$$C_{1}$$C_{2}$$C_{3}$
Figure 3: Bipartith graph representation of an MILP with three variables ($x_{1}$, $x_{2}$, and $x_{3}$), three constraints ($C_{1}$, $C_{2}$, and $C_{3}$), and AA as in Eq. (11).
@@ 5.2 SatGNN
We name SatGNN the base model proposed for all experiments. An overview of the components of our model can be seen in the diagram of Figure 4, and a summary of the symbols used is presented in C.
SatGNN is designed to take as input instances of the ONTS problem. Each instance is a mathematical program with a varying number of variables and constraints (i.e., varying number of jobs). Thus we expect our model to handle inputs of different sizes and to be invariant to the symmetries that mathematical programs exhibit (e.g., permutations of variables, constraints). For that reason, we follow the approach described in Section 4 of embedding each instance as a bipartite graph, and design SatGNN around a GNN (hence its name).
To fully represent each instance, the input graph is augmented with feature vectors for every node. Let $G=(V,E,w)$ be the graph representation of an instance $I\in\mathcal{I}$, as described in Sec. 4.1, with $V=V_{\textrm{var}}\cup V_{\textrm{con}}$ being the set of nodes and $w:E\longrightarrow\mathbb{R}$ the edge-weight function based on the variable coefficients in the constraints. Let $\boldsymbol{f}_{v}$ denote the feature vector associated with node $v$. We add four features to constraint nodes and six to variable nodes, as described in Table 2. We highlight that the features overdetermine the instance; that is, more features than necessary are provided to the model. This is common in deep learning practice and has been observed to yield positive results in combinatorial optimization applications (Nair et al., 2020; Khalil et al., 2022).
@@
${\rm NN}_{\rm var}:\mathbb{R}^{6}$ $\longrightarrow\mathbb{R}^{d}_{+}$
@@
${\rm NN}_{\rm con}:\mathbb{R}^{4}$ $\longrightarrow\mathbb{R}^{d}_{+}$
@@
Figure 4: SatGNN architecture overview, where $G$ is the bipartite graph representation of an MILP problem instance, $H$ variables represent the sets of hidden features of the nodes, and $\rm{NN}_{\cdot}$ are neural networks. The connection of $G$ and the GraphConvGraphConv operators represents both the weights of the edges and the neighborhood information.
@@ 5.3.1 Feasibility Classification
Given an instance $I\in\mathcal{I}$ of the ONTS problem and a candidate solution $\boldsymbol{\hat{z}}$, we aim for the model to classify whether $\hat{\boldsymbol{z}}$ is a feasible solution. We address this task using SatGNN by embedding the candidate solution as an additional feature in each variable node, thereby adding one extra dimension to every $\boldsymbol{f}_{v_{\text{var}}}$. Also, the model’s output is a binary value (0 or 1), indicating whether the assignment is infeasible or feasible, respectively.
As briefly described in Section 5.2, the feasibility classification is a graph classification task. It is therefore natural to take advantage of maximum likelihood training. To this end, we define $y=p(\hat{\boldsymbol{z}}\in Z|I)\in\{0,1\}$ as the target for each assignment and instance. Note that $p(\hat{\boldsymbol{z}}\in Z|I)$, which represents the probability that assignment $\hat{\boldsymbol{z}}$ is feasible, matches the label, once there is no stochastic component in the feasibility. We then use SatGNN to obtain an estimate $\hat{p}(\hat{\boldsymbol{z}}\in Z\mid I)\in[0,1]$ of the true probability. In particular, we do so by aggregating the output of the last fully-connected module ${\rm NN}_{\rm out}$, and using the sigmoid function.
The model is trained to minimize the binary cross-entropy (BCE) between the predicted probability $\hat{p}(\hat{\boldsymbol{z}}\in Z|I)\in[0,1]$ and the true probability $p(\hat{\boldsymbol{z}}\in Z|I)\in[0,1]$. We note that the minimization is performed over the parameters of the graph convolutions, as well as the parameters of ${\rm NN}_{\rm con}$, ${\rm NN}_{\rm var}$, and ${\rm NN}_{\rm out}$.
The training data is built with instances from $\mathcal{D}$ (see Sec. 5.1). However, in each $(I,Z^{\star})\in\mathcal{D}$, we only have high-quality feasible solutions in $Z^{\star}$, which is clearly not representative of the input space for the feasibility classification task. In other words, we need both feasible and infeasible candidate solutions for the model to learn the classification.
For that end, we augment $Z^{\star}$ with two other sets of candidate solutions. The first, which we call $Z_{R}$, is constructed from candidate solutions uniformly sampled from $\{0,1\}^{2JT}$, forming a representative sample of the entire domain. The second set, $Z_{N}$, is constructed from candidate solutions in the neighborhood of $Z^{\star}$. Our intuition behind $Z_{N}$ is to increase the information density of our sample near the borders of the feasible region, which is where our target function is highly nonlinear. To achieve this, we build $Z_{N}$ by randomly flipping a limited number of assignments in randomly selected candidate solutions from $Z^{\star}$.
Finally, our dataset for feasibility classification can be described as
$\mathcal{D}_{\rm feas}=\left\{(I,\hat{\boldsymbol{z}},y)\in\mathcal{I}\times Z\times\{0,1\}:(I,Z^{\star})\in\mathcal{D},\hat{\boldsymbol{z}}\in Z^{\star}\cup Z_{R}\cup Z_{N},y=p(\hat{\boldsymbol{z}}\in Z|I)\right\},$
where $p(\hat{\boldsymbol{z}}\in Z|I)$ is computed for each candidate solution [^3] Except for those in $Z^{\star}$, which are always feasible. by evaluating the constraints of $I$.
@@ 5.3.2 Optimal Solution Prediction
We want SatGNN to predict an optimal solution given an instance of the ONTS problem. We expect this task to be more challenging than the previous one, as finding an optimal solution is significantly harder than checking feasibility. Furthermore, an instance may have multiple optimal solutions, so the target for the model is not trivially well-defined.
We formulate this task through the variable bias at optimality, that is, the probability that each binary variable takes a positive value in an optimal solution chosen at random under a uniform distribution. We denote this probability vector as
$\boldsymbol{p}(\boldsymbol{z}^{\star}=1|I)=\begin{bmatrix}p(z^{\star}_{1,1}=1|I)\\ \vdots\\ p(z^{\star}_{J,T}=1|I)]\end{bmatrix}$,
where $p(z^{\star}_{j,t}=1|I)$ is the probability that the variable $z^{\star}_{j,t}$ takes a positive value in an optimal solution to instance $I$. We use SatGNN to obtain an estimate $\hat{\boldsymbol{p}}(\boldsymbol{z}^{\star}=1|I)$, which we can sample to generate candidate solutions for $I$. In particular, we use thresholding to obtain a candidate solution.
The standard classification approach for this task is to train the model to minimize the BCE with respect to the best-known solution, which serves as the target. Note that the best-known solution is indeed the target when the model has a single optimal solution and the instance has been solved to optimality. For training with the best solution, we build the dataset by selecting the best $\boldsymbol{z}^{\star}\in Z^{\star}$ for each $(I,\boldsymbol{z}^{\star})\in\mathcal{D}$, i.e.,
$\mathcal{D}_{\text{opt-b}}=\left\{(I,\boldsymbol{z}^{\star})\in\mathcal{I}\times Z:(I,Z^{\star})\in\mathcal{D},\boldsymbol{z}^{\star}=\arg\max_{\hat{\boldsymbol{z}}\in Z^{\star}}{\rm QoS}(\hat{\boldsymbol{z}};\boldsymbol{u}_{I})\right\},$
@@
Another training approach is the one proposed by Nair et al. (2020), in which multiple known solutions are used as targets, being weighted by their objective value. Note that the best-known solution is also used in this approach, alongside sub-optimal solutions. Intuitively, this multi-target approach guides the model toward the region of high-quality solutions, rather than solely toward the best-known solution, which may result in more feasible predictions.
The dataset for the multi-target approach simply requires computing the weights for each $\hat{\boldsymbol{z}}\in Z^{\star}$ of each instance in $\mathcal{D}$. More specifically, we build
$\mathcal{D}_{\text{opt-m}}=\left\{(I,\hat{\boldsymbol{z}},w_{\hat{\boldsymbol{z}}})\in\mathcal{I}\times Z:(I,Z^{\star})\in\mathcal{D},\hat{\boldsymbol{z}}\in Z^{\star},w_{\hat{\boldsymbol{z}}}=\frac{\exp({\rm QoS}(\hat{\boldsymbol{z}};\boldsymbol{u}_{I}))}{\sum_{\boldsymbol{z}\in Z^{\star}}\exp({\rm QoS}(\boldsymbol{z};\boldsymbol{u}_{I}))}\right\}.$
@@ 5.4 SatGNN-based Heuristic
Although we could use the SatGNN by itself as a heuristic for the ONTS problem by sampling its output to generate candidate solutions, it is unlikely that these solutions will be useful. Because our problem has constraints, an error in a single variable may result in an infeasible candidate. Given that ONTS problems of moderate size already have thousands of variables [^4] The smallest instance evaluated by Rigo et al. (2022) has 1746 binary variables, considering both $x$ and $\phi$ variables., infeasibility is likely to occur. To circumvent this limitation, we use the deep learning model in conjunction with a mathematical programming solver as a primal heuristic. In particular, we propose three approaches for doing so, which instantiate recent GNN-guided baselines (Nair et al., 2020; Han et al., 2023).
@@ 5.4.1 Warm-starting
A straightforward way to aid most commercial-grade solvers is to warm-start them with good candidate solutions. For example, the SCIP solver (Bestuzheva et al., 2021) accepts complete and partial solutions to the problem, which are used to guide the inner heuristics during the optimization process. We use the output of the model to determine which variables will compose the partial solution provided to the solver based on the confidence of the model’s prediction. In other words, we warm-start the solver with partial solutions built using only the assignments for variables for which the model predicts values closest to 0 or 1.
Formally, let $\hat{\boldsymbol{p}}(\boldsymbol{z}^{\star}=1|I)$ be the output of SatGNN given an instance $I\in\mathcal{I}$ as defined in Sec. 5.3.2. Our definition of confidence for a given variable comes from the maximum predicted marginal probability. For instance, the confidence with respect to variable $x_{j,t}$ of an instance $I$ is
$\kappa(x_{j,t}|I):=\begin{cases}\hat{p}(x_{j,t}^{\star}=1|I),&\text{if}\quad\hat{p}(x_{j,t}^{\star}=1|I)\geq 0.5\\ \hat{p}(x_{j,t}^{\star}=0|I)=1-\hat{p}(x_{j,t}^{\star}=1|I),&\text{otherwise}.\end{cases}$
The confidence associated with the $\phi_{j,t}$ variables can be described similarly. Note that the confidence is always between 0.5 and 1. Furthermore, we write $\boldsymbol{\kappa}(\boldsymbol{z}^{\star}|I)$ to denote the confidence vector over all variables.
@@
Note that warm-starting does not configure a heuristic solution to the problem, as it just modifies the behavior of the heuristics already present in the algorithmic solution. Therefore, we maintain optimality and feasibility guarantees, even in the worst-case scenario when the model provides the solver with bad, infeasible solutions.